10/15/99 Day shift summary

Ashkan_Alavi-Harati (alavi@fnal.gov)
Fri, 15 Oct 1999 15:42:17 -0500

Shift Summary for day shift, 15/Oct/99

On Shift: Jason Hamm, Ashkan Alavi

Hours of beam available at all intensities: 6.5
Hours of beam available at nominal intensity: 5.3
Hours of E799 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.8

Accounting of lost time:

Accelerator planned and unplanned downs, tail of the laser scan, stop and
starting runs + drt_fresh_start.
========================================================================
Runs ended this shift:

Run 14843 - 7.9 M evts - BEAM_799
Run 14844 - LASER_SCAN
Run 14845 - 5.3 M evts - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

UP*76-78, UPS026
========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

Relatively smooth 799 datataking. Also did a laser scan.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14843 143 07:00 Take over the ongoing run

08:30 onsplit tape UPK010 mounted of ktev1

08:50 onsplit tape UPF020 mounted of ktev2

230 09:00 B4 frigde is toggling, est. 10-15 min for beam

234 09:03 Beam is back earlier than estimated time.

278 10:00 MCR turns our beam off for about 30 minutes
extraction studies. Stop the run.

14844 10:11 Start Laser scan.

10:15 Flip the analysis magnet to positive.

10:41 onsplit tape UPT006 mounted of ktev4.

10:48 Beam returns with low intensity but we are not
finished with the laser scan yet.

10:55 laser scan ends.

14845 1 11:05 Start the run, but beam is gone due to linac
problems. est. 10 min.

10 11:20 beam returns.

15 11:25 back to the nominal intensity.

42 12:03 beam is gone, left bends fridge permit toggle.

49 12:13 beam returns.

67 12:37 Liquid power supply problems, est 10 min.

12:50 onsplit tape UPM017 mounted of ktev1.

101 13:22 beam returns.

13:15 onsplit tape UPH011 mounted of ktev2

14:15 dart people show up to work on the damp problem.

14:35 beam structure is crappy, Jason calls MCR, people
are playing around at the switchyard.

168 14:50 Stop the run to practice drt_fresh-start on
Margaret's request. They may have broken our nodes
so the drt_fresh_start may not work.

14846 1 15:10 drt_fresh_start. Start the run. See the note below
from the dart people.

11 15:26 ktevana errors, KFREND.
========================================================================
Follow up Notes on Previous Shift

Did drt_fresh_start after stopping 14843.

Did the laser scan.
========================================================================
Items to Follow up on Next Shift

They think the dms server is not sending out commands. So the problem is
not the damp server.
If the problem occurs again, please run:

ktev3> reaper

(or the command may be 'reaper ktev')

and also page the dart primary (via the helpdesk) so they can look at it
live, if we can spare the time for them to invesitage it.

(see logbook p102 for full description)